OpenStudy (anonymous):

Simplify the quantity negative 3 times a times b to the negative third times c to the fifth all to the negative 2

OpenStudy (anonymous):

\[(-3ab^{-1/3}c^{1/5})^{-2}\] is that it?

OpenStudy (anonymous):

the question i mean

OpenStudy (anonymous):

(-3ab^-3c^5)^-2

OpenStudy (anonymous):

\[(-3ab ^{-3}c ^{5})^{-2} = 1/(-3ab ^{-3}c ^{5})^{2} = 1/(9a^2b^{-6}c^{10})\]

OpenStudy (anonymous):

you can also write: \[b^{6}/(9a^{2}c^{10})\] because the negative exponent is on the denominator.

OpenStudy (anonymous):

Great job! that was the form I was looking for, thank you